XxSavageForLifexX:

b squared+18b+81 factoring polynomials plzzz gcf first then factor

jhonyy9:

b^2 +18b +81 = ? do you know formula (a+b)^2 = ?

jhonyy9:

do you can getting two numbers with sum equal 18 and product 81 ?

XxSavageForLifexX:

no i dont

jhonyy9:

one example : a^2 +4a +4 = ? can you factorizing it ? you need getting two numbers with sum equal 4 and product equal 4 can you get ?

jhonyy9:

so 2 and 2 bc. 2+2=4 and 2*2=4 than a^2 +4a +4 = (a+2)^2 do you understand it now ?

jhonyy9:

courage this is easy

jhonyy9:

how can you make 18 from 2 numbers assumed ?

jhonyy9:

for example 6+12 =18 but 6*12 = 72 and you need getting product equal 81

jhonyy9:

any idea you dont have ? i m here to help you

jhonyy9:

so 8 and 10 will be right ?

XxSavageForLifexX:

ye it would thnx

jhonyy9:

no bc. 8+10 = 18 but 8*10 = 80 and you need product equal 81

XxSavageForLifexX:

ik i just gave up dudei hate this i hate all math

jhonyy9:

dont give up bc. we here wan helping you but you need collaborate

XxSavageForLifexX:

its difficult

jhonyy9:

not is looke please than you know that a+a=18 and a*a=81 so how you get this number of a ?

XxSavageForLifexX:

idk im feeling dumb

jhonyy9:

looke please a+a=18 => 2a = 18 divide both sides by 2 so what will get a = ?

XxSavageForLifexX:

9?

jhonyy9:

exactly so 9+9 = ?

XxSavageForLifexX:

18

jhonyy9:

very nice and 9*9 = ?

XxSavageForLifexX:

81

jhonyy9:

yes sure so now your polinom how will be factorized ?

jhonyy9:

looke please on the top what i ve wrote there

XxSavageForLifexX:

this is where its confusing me tho

jhonyy9:

ok. than check this please bc. you know now that 9+9=18 and 9*9=81 you can rewriting your polinom in this form b^2 +18b +81 = b*b +9b +9b +9*9 = ? what can you factorizing out from these terms gruping them ?

XxSavageForLifexX:

19b+81...? idk

jhonyy9:

ok so from first two terms you factoriz out the b and from 3rd and 4th the 9 and will get = b(b+9) + 9(b+9) = and now what is common factor there ?

XxSavageForLifexX:

b^2+9b+9b+81

jhonyy9:

wait please can you answer my question ? what is common factor there b(b+9) +9(b+9) =

XxSavageForLifexX:

I DONT KNOW

jhonyy9:

(b+9) is common there and than you factoriz it out what will get ?

XxSavageForLifexX:

why cant u see i dont get this

jhonyy9:

you get (b+9)(b+9) = ?

XxSavageForLifexX:

no i dont this is what i get for being homeschooled until high school

jhonyy9:

than you know that a*a = a^2 so than (b+9)(b+9) = ?

jhonyy9:

@ThisGirlPretty do you can explain it more understandably please ?

jhonyy9:

ok. so than note b+9= x so in this case (b+9)(b+9) = x*x = ?

jhonyy9:

i ve wrote above that a*a = a^2 so than x*x = ?

ThisGirlPretty:

Oh wait are you suppose to do 9 x 9 = ?

jhonyy9:

no given b^2 +18b +81 and this polinimial you need factoriz it getting the gfc

jhonyy9:

sorry gcf

Hero:

Is @XxSavageForLifexX online right now?

jhonyy9:

yes was here now one moment

jhonyy9:

but than you check he now ve go away

jhonyy9:

@ThisGirlPretty ty all your contribution ,collaboration

ThisGirlPretty:

Anytime ! :)

jhonyy9:

i saing same any time i m here to help you than you like ,wan

ThisGirlPretty:

Yes sir ^-^ !

XxSavageForLifexX:

i still dont understand this